OpenStudy (anonymous):

Rene is going to the lake to visit some friends. If the lake is 60 miles away, and Rene is driving at 40 miles per hour the entire time, how long will it take her to get to the lake?

OpenStudy (anonymous):

an hour and a half

OpenStudy (anonymous):

1 hour is 40 miles. 60-40= 20. 20=1/2 of 40, 1/2 of one hour is 30 minutes.

OpenStudy (anonymous):

or if your doing a unit rate. 40/1 = 60/x. 60/40*1 = x = 1.5